OpenStudy (anonymous):

At a restaurant, the bill comes to $55. You decide to leave a 19% tip. How much did you leave for the tip?

OpenStudy (anonymous):

@wio @ganeshie8

OpenStudy (anonymous):

Multiply $55 x .19

OpenStudy (apricot):

$10.45 is your answer.

jimthompson5910 (jim_thompson5910):

You can do a bit of rounding 19% is close to 20%, which is 1/5 55*(1/5) = 55/5 = 11, which is close to $10.45

OpenStudy (anonymous):

thanks everyone!